OpenStudy (anonymous):

Can you help me with problems "d" and "e" of the following trig app question?

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

For (d), solve sin[0.2094(t-30)]=sin 270. Hmm.. I'm not sure

OpenStudy (callisto):

Since, sin 270 is the minimum for a normal sine curve. So, equate the sine part in the equaton and sin270. I think you'll get something

OpenStudy (anonymous):

Does this work? Set 3π/2=(.2094)t t=22.5 minutes or 22.5+(n)(30) where n is a positive integer.

OpenStudy (callisto):

[0.2094(t-30)]= 3pi /2 You have a horizontal shift

OpenStudy (anonymous):

Does that mean its wrong?

OpenStudy (callisto):

I think.. [0.2094(t-30)]= 3pi /2 would be more appropriate.. Not sure

OpenStudy (anonymous):

Wait so the answer is 3pi/2?

OpenStudy (callisto):

Nope, that is the equation only you need to solve t

OpenStudy (anonymous):

I dont uderstand though, [0.2094(t-30)] equals to 3pi /2. What I am to solve then?

OpenStudy (callisto):